OpenStudy (anonymous):

Find the difference of (−5x2 − 4x3 − 3) − (9x3 + 3x − 8x2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2.

OpenStudy (anonymous):

(-25x - 64x - 3) - (729x + 3x - 64x) ????

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

you need to subtract like terms x^2 minus x^2

OpenStudy (anonymous):

so the -5x^2 and -8x^2?

OpenStudy (anonymous):

yes, and you get?

OpenStudy (anonymous):

add 8 to boith sides and get 3x^2

OpenStudy (anonymous):

yes your answer is right but your method is fishy.. any go on :) i would say -5x^2--8x^2 = -5x^2+8x^2=3x^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do the same for the rest, and give me your answer :)

OpenStudy (anonymous):

-4x^3 and 9x^3

OpenStudy (anonymous):

yes and ur answer?

OpenStudy (anonymous):

or should the 9 be negative because the minus sign in between the two equations??

OpenStudy (anonymous):

well, the 9 is postive but you have to subtract it from 4..

OpenStudy (anonymous):

-4x^3-9x^3 is the same as -4x^3-+9x^3

OpenStudy (anonymous):

-13x^3?

OpenStudy (anonymous):

yes, and the last two terms, do they have partners?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol, correct..not the same terms.. so your answer would be all the stuff together.. i'll give you a start: 3x^2-13x^3and the last two terms without partners

OpenStudy (anonymous):

so 3x^2 - 13^3 + 3x - 3?

OpenStudy (anonymous):

yes :) well done, by the way.. there is a difference between (5x)^2 and 5x^2.. do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(5x)^2= 5^2x^2=25x^2 5x^2 is just 5x^2 :) next time you should write your question with the exponents to avoid complications :)

OpenStudy (anonymous):

ok thank you for all the help

OpenStudy (anonymous):

you are welcome, you are a great learner :)